OpenStudy (anonymous):

let f(x)=x/(x-2) . find a function y=g(x) so that (f°g)(x)=3x g(x)= ?? anybody can help??

OpenStudy (anonymous):

there is a catch here, x can't be 3. but otherwise the solution is easy

OpenStudy (anonymous):

actually i don't understand the whole thing

OpenStudy (anonymous):

let \[g(x) = y\] so, \[f(g(x)) = 3x \Rightarrow f(y) = 3x \Rightarrow \frac{y}{y-2} = 3x\] which gives \[y = \frac{6x}{3x-1}\] Hence \[g(x) = \frac{6x}{3x -1 }\] for a check, just plug in the f(g(x)) and see it equals 3x, but remember \[x \neq \frac{1}{3}\]

OpenStudy (anonymous):

yes, sure !! thank you